# Forging GEMs: Advancing Greek NLP through Quality-Based Corpus Curation and Specialized Pre-training

### v1

@@ A.1 Procedure
The computational requirements varied significantly based on architectural complexity. For instance, models with efficient pre-training objectives like ELECTRA were among the lightest, while architectures designed for long contexts, such as Longformer, were the most computationally intensive.
A.2 Computational Requirements
Table 10 summarizes the computational requirements for pretraining the model architectures evaluated in this work. Training times varied significantly based on model complexity and context length, with ELECTRA being the most efficient and LongFormer the most computationally intensive due to its extended context window. Due to resource and time constraints, not all architectural variants were trained on all corpora; we report training times only for the configurations that were actually pretrained.
@@ A.2 Computational Requirements
The reported training times are highly dependent on the selected batch size for each experiment and thus should be interpreted in relation to the corresponding configuration.

### v2

@@ Appendix A Appendix
Table 9: Pre-training Hyperparameters and Configuration.
NER MCLTC (Volume) MCLTC (Chapter) MCLTC (Subject) NLI
lr dropout lr dropout lr dropout lr dropout lr dropout
Greek-BERT 3e-05 0.1 2e-05 0.1 2e-05 0.2 2e-05 0.2 2e-05 0.0
Greek-Legal BERT 5e-05 0.2 2e-05 0.2 2e-05 0.1 2e-05 0.0 3e-05 0.2
Greek-Legal RoBERTa 2e-05 0.1 2e-05 0.1 2e-05 0.2 3e-05 0.2 2e-05 0.0
Greek-Media LongFormer 2e-05 0.1 2e-05 0.1 2e-05 0.1 3e-05 0.2 5e-05 0.0
GEM-RoBERTa 5e-05 0.2 2e-05 0.1 2e-05 0.1 5e-05 0.2 3e-05 0.2
GEM-ELECTRA 3e-05 0.2 2e-05 0.1 2e-05 0.1 2e-05 0.0 3e-05 0.2
GEM-ConvBERT 3e-05 0.2 2e-05 0.1 2e-05 0.1 2e-05 0.1 2e-05 0.0
GEM-ModernBERT 5e-05 0.1 2e-05 0.1 3e-05 0.0 2e-05 0.2 2e-05 0.2
GEM-LongFormer 2e-05 0.2 2e-05 0.2 2e-05 0.2 2e-05 0.3 4e-05 0.1
GEM-RoBERTa 2e-05 0.0 2e-05 0.0 2e-05 0.1 2e-05 0.0 2e-05 0.0
GEM-XLM-RoBERTa 2e-05 0.0 2e-05 0.1 2e-05 0.1 2e-05 0.0 2e-05 0.1
Table 10: Best hyperparameters for finetuning the models in our experiments.
The computational requirements varied significantly based on architectural complexity. For instance, models with efficient pre-training objectives like ELECTRA were among the lightest, while architectures designed for long contexts, such as Longformer, were the most computationally intensive. Table 11 summarizes the computational requirements for pretraining the model architectures evaluated in this work. Training times varied significantly based on model complexity and context length, with ELECTRA being the most efficient and LongFormer the most computationally intensive due to its extended context window. Due to resource and time constraints, not all architectural variants were trained on all corpora; we report training times only for the configurations that were actually pretrained.